You have just learned that B&B has undertaken a major expansion that will change its expected free cash flows to−$10 million in 1 year, $20 million in 2 years, and $35 million in 3 years. After 3 years, free cash flow will grow at a rate of 5%. No new debt or preferred stock were added, the investment was financed by equity from the owners. Assume the WACC is unchanged at 11% and that there are still 10 million shares of stock outstanding. What is its horizon value (i.e., its value of operations at year three)? What is its current value of operations (i.e., at time zero)?
Answer #1

What is its horizon value (i.e., its value of operations at year three)

=(free cash flow at year 3*(1+growth rate))/(WACC-growth rate)

=(35*(1+5%))/(11%-5%)

=612.50 million

What is its current value of operations (i.e., at time zero)

=-10/(1+11%)^1+20/(1+11%)^2+35/(1+11%)^3+((35*(1+5%))/(11%-5%))/(1+11%)^3

=480.67 million

value per share=480.67/10=48.07 per share

the above is answer..